OpenStudy (jjuden):

FAN AND MEDAL Jeremiah is factoring 0.64t^24 – 9 by using the rule a^2 – b^2 = (a + b)(a – b). What will he use for the value of a?

OpenStudy (jjuden):

A 0.08t^12 B 0.8t^12 C 0.64 D 0.8

OpenStudy (aravindg):

So basically 0.64t^24 – 9=a^2-b^2=(a+b)(a-b) Which number squared gives 0.64t^24 Compare and get your answer.

OpenStudy (jjuden):

a

OpenStudy (aravindg):

Nope 0.08*0.08=.0064